Find the maximum and minimum values of the function f(x,y)=2x2+3y2−4x−5 on the domain x2+y2≤324.

The maximum value of f(x,y). and List the point(s) where the function attains its maximum as an ordered pair, such as (-6,3), or a list of ordered pairs if there is more than one point, such as (1,3), (-4,7).
1
Expert's answer
2020-04-08T16:39:22-0400

"f'_x=4x-4,f_y'=6y" then the stationary point is "(1,0)."

"f''_{x^2}=4,f''_{xy}=0,f''_{y^2}=6" then "(1,0)" is point of the minimum. "f(1,0)=-7"

Examine the "f(x,y)" at the boundaries of the region

"L(x,y,\lambda)=2x^2+3y^2-4x-5+\lambda(x^2+y^2-324)"

"\lambda'_x=4x-4-2\lambda x,\lambda'_y=6y+2\lambda y" if "\lambda'_x=\lambda'_y=0" and "x^2+y^2=324" then

"\lambda=-3,x=-2,y=8\sqrt5,y=-8\sqrt5"

"\lambda''_{x^2}=4+2\lambda,\lambda''_{xy}=0,\lambda''_{y^2}=6+2\lambda" hence "d^2\lambda<0" hence "(-2,8\sqrt5),(-2,-8\sqrt5)" is points of the conditional maximum. Then "f(-2,8\sqrt5)=f(-2,-8\sqrt5)=971"